These were accepted as a reasonable account of the last meeting.
Script has been created and will be in the stable release w/c 27th March. There have been 16 reported so far with the largest time difference being 5 seconds. Kerberos will cope with a difference up to 2 minutes. More responses expected once in stable release.
Neil spoke to Jon and he's keen to have a chat. He would like bayes.ed for publicity reasons but but a redirect would be possible. Also might require some 'bayes' email addresses.
All pages with review date prior to August 2015 should be reviewed by the end of April.
blog.inf is now firing extra logging and it seems to be working.
All 500 tapes will be handed over to the new owner but must be wiped first. Neil's estimate is that it will take 2 weeks. We'll see .....
Blog articles completed:
Blog articles suggested:
Report from Computing Executive Group
The sickness absence reporting policy was discussed with no objections.
Reports from Units
These are now being sent to all COs so please look at and fix any problems caused by your machines!
Agreed that this is no longer required.
We discussed whether multicast DNS was still required. It was suggested that it would only be useful on desktops and perhaps required for labs. It could, however, be removed from network servers.
Alastair stressed that this is bogus when seen at boot time but if seen at any other time, it does indicate a problem.
US's wiki page is here.
Topics for discussion
Ian noted that both Lantronix and IPMI consoles have historically been fairly reliable - so what, if anything, has recently made IPMI less reliable?
ipmitooldirectly, rather than via
conserver. Ian noted that our problems persist when a direct connection via
ipmitoolis used - this has been tested.
The set up at TCU sounded very similar to IS with little/no integration between groups. One point which we might want to consider is their policy of allocating a member of the support team to new staff members for their first week to help answer their IT queries and give a general introduction to the services provided.
Neil is awaiting a response from David Bell regarding furniture and will keep us posted. As far as we know, things are still on schedule but there is a 'return to AT' meeting next week where Martin will update on progress.
Craig would like suggestions for next week's TDM. Failing that, we can all get on with SL7 upgrades.
The next meeting will be on Wednesday 12th April at 10:00, in IF-4.31, chaired by Craig.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh